# On the long-time limit of the mean curvature flow in closed manifolds

### v1

@@ Acknowledgments
A.M. thanks Alec Payne and Felix Schulze for their interest and helpful comments. In the course of preparing the article he was supported by CPH-GEOTOP-DNRF151 from the Danish National Research Foundation and CF21-0680 from the Carlsberg Foundation (via GeoTop and N.M. Møller respectively) and is grateful for their assistance. A.S. is supported by the AMS-Simons travel grant. He is grateful to the Copenhagen Centre for Geometry & Topology (GeoTop) at the University of Copenhagen, especially N.M. Møller, for the invitation to visit in Spring 2024 and participate in the meeting "Masterclass: Recent Progress on Singularity Analysis and Applications of the Mean Curvature Flow". Part of the work was initiated during the visit.
@@ 2. Background and preliminary lemmas
Let $N^{n+1}$ be a Riemannian manifold and $X:M\to N^{n+1}$ be an embedding of a manifold $M^{n}$ realizing it as a $2$–sided smooth closed hypersurface of $N$, whose image by abuse of notation we also refer to as $M$. Then the mean curvature flow $M_{t}$ of $M$ is given by (the image of) $X:M\times[0,T)\to N^{n+1}$ satisfying the following:
@@
[…]
@@
  • The Brakke flow
  • The weak set flow and level set flow
  • The surgery flow
The Brakke flow, introduced by Brakke in his thesis [5] is a geometric measure-theoretic definition in terms of (typically, integral) varifolds that satisfy the eponymous Brakke inequality. Here we present a definition essentially due to Ilmanen [29], and the expository is from [4]. An $n$-dimensional Brakke flow defined on the time-interval $I\subset{\mathbb{R}}$ in a $(n+1)$-dimensional manifold $(N,g)$ is a family of Radon measures $(\mu_{t})_{t\in I}$ such that:
  • For a.e. $t\in I$ the measure is integer $\mathcal{H}^{n}$-rectifiable and the associated varifold has locally bounded first variation with variational vector $\vec{H}$ in $L^{1}$.
  • For any compact set $K\subset N$, and any $[t_{1},t_{2}]\subset I$,
[…]
  • For any $[t_{1},t_{2}]\subset I$ and $u\in C_{c}^{1}(N\times[t_{1},t_{2}])$,
[…]
One can quickly see that the Radon measures associated to a smooth/classical mean curvature flow, that is a family of smooth manifolds satisfying (2.1), is a Brakke flow. A Brakke flow suffers from a serious problem: as (2.2) is only an inequality, the whole connected components of (the support of) a Brakke flow are allowed to disappear instantaneously! In the same monograph, Ilmanen [29] (see also White [53]) gives a construction via elliptic regularization of a "well-behaved" Brakke flow out of smooth initial data for which these issues don’t occur. Such a Brakke flow is known to be:
@@
Even amongst such well-behaved Brakke flows there happens to be an issue of uniqueness, which brings us to the level set flow which originated from the viscosity method [18, 10]. Later Ilmanen [28] and White [50] introduced a purely set-theoretic definition using the avoidance principle. The classical avoidance principle says that two disjoint hypersurfaces in a manifold $N$ will stay disjoint, at least as long as one of them is compact. The weak set flow is defined to be a closed subset $\mathcal{M}$ of $N\times{\mathbb{R}}$ such that if a smooth mean curvature flow does not intersect $\mathcal{M}\cap\{t\}$, then it is disjoint from $\mathcal{M}\cap\{t^{\prime}\}$ for all $t^{\prime}>t$. White [55] proved the following avoidance principle for weak set flows.
@@
Indeed the support of a Brakke flow is a weak set flow, as shown in Ilmanen’s monograph. Although weak set flows may not be unique starting from a given hypersurface, there is a canonical weak set flow associated to any initial data called the level set flow: for a given set $A$ is as the envelope of all families of sets initially agreeing with $A$ that satisfy the avoidance principle. As an aside we point out that in fact the level set flow can even be defined this way for closed initial data that is not necessarily a regular hypersurface. While the existence and uniqueness are guaranteed, the level set flow of initial data may develop interior even if it is smooth, as first noted by Ilmanen and White [52] (and recently appeared in [31], see also related works [1, 34, 32]) – this fattening can be seen as an indication of nonuniqueness for level set flows. To study the level set flow even in the presence of fattening Hershkovits and White [26] introduced the notions of innermost/outermost flows. Given a closed hypersurface $\Sigma\subset{\mathbb{R}}^{3}$ that is the boundary of a domain $\Omega$, we can define:
@@
It is not immediate to define inner and outermost flows in general $3$–manifolds from a $2$–sided surface $M$ because such a surface need not be separating, as a cross section of $S^{2}\times S^{1}$ illustrates. In the sequel we will define such flows as the following: for a choice of normal vector $\nu$ consider the domains $\mathcal{M}^{-}$ and $\mathcal{M}^{+}$ bounded by $M$ and $M\pm\epsilon\nu$ respectively, for some small $\epsilon>0$, where $M\pm\epsilon\nu$ are surfaces that are generated by pushing $M$ in $\epsilon\nu$ direction by the exponential map. Then we define the inner and outermost flows to be the boundary component of the level set flows of $\mathcal{M}^{-}$, $\mathcal{M}^{+}$ respectively emanating from $M$.
The third weak flow is the mean curvature flow with surgery, where one preemptively cuts out regions developing high curvature, classifies them topologically, and then continues the flow in a piecewise manner, see [23, 33, 2, 24, 15]. Locally, the high curvature regions will be modeled on round spheres or cylinders in coordinate patches which will bound solid balls/cylinders respectively. Regions where high curvature transitions into low curvature (i.e. "neckpinches") the surface are modeled on cylinders and at these points one "cuts" the neck at an appropriate spot and places in caps/discs of controlled geometry, in particular so that the remaining low curvature regions have curvature upper bounded by a controlled constant (which is amongst the analytical difficulties in establishing the flow with surgery). We note that the high curvature regions after cap placement will be homeomorphic to either $S^{n}$ or $S^{n-1}\times S^{1}$, which makes the flow with surgery a useful tool in topology. This particular consequence won’t be so relevant in this work but the well controlled, concretely understandable change of topology across surgeries will be helpful in the sequel.
@@ 2.1. Almost regular flow and generic approximation
While these notions of weak flows seem quite different, in $3$–manifolds there is a relatively complete picture to relate all of them. In [4], Bamler and Kleiner introduced a new notion of weak flow called almost regular flow which are well behaved Brakke flows in the sense described above along with an extra integrability condition on what they call scale functions, which measure the size of neighborhoods in which the flow is smooth. We refer the readers to [4] for the definition and properties of almost regular flows. Here we only state properties of almost regular flows that we reference:
@@
    Generic flows, i.e. ones which only encounter mean convex singularities (see discussion and references below) are almost regular.
@@
    The innermost/outermost flows that start from a closed smooth embedded surface can be approximated by generic flows as a consequence are almost regular flows for $t\geq 0$. In particular, if the level set flow from a closed smooth embedded surface is nonfattening then it is almost regular.
@@
[…]
While the almost regular flows have many nice properties, they still possibly develop complicated singularities which obscures the possible topological change of the flow through them. On the other hand, as we have described before, the mean curvature flow with surgery has a clear description of the topological change of the flow. Therefore, we would like to use mean curvature flow with surgery to approximate the almost regular flows we use in the sequel, particularly in showing (some of) the applications mentioned in the introduction.
In order to use mean curvature flow with surgery to approximate the almost regular flows we use in practice, we first use a well behaved (unit regular, cyclic) Brakke flow with only cylindrical and spherical singularities to approximate them. Such flows we call generic. It has been conjectured by Huisken that mean curvature flow with generic initial data can only have cylindrical and spherical singularities, and recently there has been various progress on this conjecture, [12, 7, 8, 46, 45]. Using the result of Chodosh, Choi, Mantoulidis and Schulze [7] and Chodosh, Choi, and Schulze [8], Bamler and Kleiner [4] proved that a mean curvature flow starting from a generic closed embedded surface in a three-manifold can only have cylindrical and spherical singularities. Furthermore, Daniels-Holgate [15] proved that the mean curvature flows with surgeries can approximate a mean curvature flow with only cylindrical and spherical singularities, where the approximation is in the Hausdorff distance. In addition, this approximation is smooth away from the singular set. Combining all the ingredients above gives the following approximation result in ${\mathbb{R}}^{3}$:
@@
Note that in the classical case i.e. $N={\mathbb{R}}^{3}$ properly embedded hypersurfaces are automatically orientable. If $N$ is more generally simply connected it is not hard to see this true by an intersection number argument, but we point out it does not suffice if $N$ is merely orientable with as the example ${\mathbb{R}}P^{2}\subset{\mathbb{R}}P^{3}$ shows. With this in mind, we claim that the following holds true in general closed $3$–manifolds:
@@
Let $M$ be a closed, properly embedded $2$–sided surface $M$ in a closed $3$–manifold $N$. Then the inner and outermost flows of $M$, as defined above, can be approximated by mean curvature flows with surgeries in Hausdorff distance over any fixed time interval and the approximation is smooth away from the singular set.
@@
The main point is to establish the existence of generic mean curvature flow ala [8] and its subsequent approximation by surgery flows, then the statement about inner and outermost flows follows exactly as in the Euclidean case by an approximation argument as discussed in section 7 of [4]. As is well known, in a blowup limit singularities will still be modeled on solitons in ${\mathbb{R}}^{3}$ so we mostly just supply a sketch of this argument. Noting that because $M$ is 2-sided one may meaningfully discuss one-sided perturbations of $M$ and the point made in the previous sentence one readily sees from [7, 8] that if one encounters a nongeneric singularity of multiplicity one that the initial data can be perturbed to avoid the singularity at that spacetime point. Because $M$ is a compact surface it has finite genus and so by genus monotonicity (see Section 11 of [7]) and the classification of genus 00 shrinkers by Brendle [6], only finitely many (arbitrarily small) perturbations of the initial data are required to produce a generic flow up at least to the first time a high multiplicity tangent flow develops. Now all these generic flows, as shown in section 7 of [4], are almost regular flows; note this uses the resolution of the mean convex neighborhood conjecture by Choi, Haslhofer, and Hershkovits [11] which the authors note carries over in the curved ambient setting. Because under rescaling as in singularity analysis the ambient metric converges smoothly to the Euclidean metric, one can see by contradiction that the resolution of the multiplicity one conjecture applies in the curved ambient setting because all the relevant geometric quantities can be arranged to evolve as they do in the flat case up to arbitrarily small errors – their argument at a high level is by showing in the case of high multiplicity convergence in a tangent flow that a certain quantity satisfies contradictory bounds so this suffices for our needs. As a result we get that for generic choice of data there will be a flow along which only mean convex singularities may develop for all time; by the resolution of the mean convex neighborhood conjecture [11] and the barrier argument of Hershkovits and White [26, 27], it is unique among almost regular flows starting from $M$. The approximation theorem of Daniels-Holgate also applies in this setting; most importantly besides the mean convex neighborhood conjecture the required local estimates for surgery also hold in the general ambient setting as shown by Haslhofer and Ketover [25]. ∎
We next present some properties of such almost regular flows, where $M$ is as usual a properly embedded $2$–sided surface in a given closed $3$–manifold $N$.
@@
Suppose that $M_{t}$ is a generic or inner/outermost flow out of $M$. Then $M_{t}$ will also be 2–sided for all regular times $t$.
@@
It’s obvious if $M_{t}$ is smooth then it will remain 2–sided, because the flow is an isotopy. In the case that $M_{t}$ has singularities, we first observe that it is easy to see for the approximating smooth flows $S_{t}$ to $M_{t}$ produced in [15] that $S_{t}$ will remain orientable if $M$ is. The convergence of the approximating surgery flows is smooth during regular times, giving the claim. ∎
An important subtlety to remember though, as mentioned in the introduction, is that orientability is not necessarily inherited by the limit we discuss in the next section. In the next statement we write $M\sim\sigma$ to indicate that $M$ is homologous to $\sigma$.
@@
Again using that the approximation by surgery flows is smooth during regular times, one only needs to note that surgery does not change the $2$nd homology class of a regular surface. ∎
@@
As in the previous proofs it suffices to note this is the case for approximating surgery flows. Now in the surgery components are either deleted or cylindrical regions are cut and replaced with caps and both of these actions do not increase the genus, giving the claim. ∎
We show in the next section that an almost regular flow will either go extinct or flow to a potentially unstable smooth minimal surface, possibly with multiplicity. If the minimal surface is unstable we will afterwards wish to perturb the flow to ensure it doesn’t flow to it, but as discussed in the introduction for technical reasons it seems difficult to carry this out by just perturbing the initial data if the multiplicity is high (i.e. greater or equal than $2$). With this in mind we will consider the following piecewise almost regular flows in the sequel:
@@
Let $M$ be a properly embedded, $2$–sided surface of a $3$–manifold $N$. Then a piecewise almost regular flow $M_{t}$ emanating from $M$ defined on $[0,\infty)$ is given by an increasing (possibly finite) sequence of times $t_{i}\to\infty$ for which
@@
    $M_{t}$ on $[0,t_{1}]$ is an almost regular flow out of $M$ satisfying the conclusions of Proposition 2.2 above and so that $M_{t_{1}}$ is a smooth surface.
@@
    Denoting by $M_{t}^{1}$ on $[0,t_{1})$ the flow from above, we let $\tilde{M}^{1}_{t_{1}}$ be a surface that is isotopic to $M^{1}_{t_{1}}$, and we define $M^{2}_{t}$ out of this perturbed surface $\tilde{M}^{1}_{t_{1}}$ as above.

### v2

@@ Acknowledgments
A.M. thanks Alec Payne and Felix Schulze for their interest and helpful comments. In the course of preparing the article he was supported by CPH-GEOTOP-DNRF151 from the Danish National Research Foundation and CF21-0680 from the Carlsberg Foundation (via GeoTop and N.M. Møller respectively) and is grateful for their assistance. A.S. is supported by the AMS-Simons travel grant. He is grateful to the Copenhagen Centre for Geometry & Topology (GeoTop) at the University of Copenhagen, especially N.M. Møller, for the invitation to visit in Spring 2024 and participate in the meeting "Masterclass: Recent Progress on Singularity Analysis and Applications of the Mean Curvature Flow;" part of this work was initiated during the visit. The authors also thank the anonymous referees for their careful reading and helpful feedback, which helped improve the exposition of the article.
@@ 2. Background and preliminary lemmas
Let $N^{n+1}$ be a Riemannian manifold and $X:M\to N^{n+1}$ be an embedding of a manifold $M^{n}$ realizing it as a 22–sided smooth closed hypersurface of $N$, whose image by abuse of notation we also refer to as $M$. Then the mean curvature flow $M_{t}$ of $M$ is given by (the image of) $X:M\times[0,T)\to N^{n+1}$ satisfying the following:
@@
An immediate consequence of the first variation formula and the evolution equation above is that the mean curvature flow is the gradient flow of the area functional, in the sense that $\frac{d}{dt}\text{Area}(M_{t})=-\int_{M_{t}}|\vec{H}|^{2}d\mu_{t}$, where $\mu_{t}$ is the area measure; from this one sees the mean curvature is well adapted to finding minimal surfaces, which are critical points of the area functional.
@@
  • The Brakke flow.
  • The weak set flow and level set flow.
  • The flow with surgery.
The Brakke flow, introduced by Brakke in his thesis [Brakke78], is a geometric measure-theoretic definition in terms of (typically, integral) varifolds that satisfy the eponymous Brakke inequality. Here we present a definition essentially due to Ilmanen [Ilmanen94_EllipReg], and the expository is from [BamlerKleiner23_Multiplicity1]. An $n$-dimensional Brakke flow defined on the time-interval $I\subset{\mathbb{R}}$ in a $(n+1)$-dimensional manifold $(N,g)$ is a family of Radon measures $(\mu_{t})_{t\in I}$ such that:
  • For a.e. $t\in I$, the measure is integer $\mathcal{H}^{n}$-rectifiable and the associated varifold has locally bounded first variation with variational vector $\vec{H}$ in $L^{1}$.
  • For any compact set $K\subset N$, and any $t_{1},t_{2}]\subset I$,
    $\int_{t_{1}}^{t_{2}}\int_{K}|\vec{H}|^{2}d\mu_{t}dt<+\infty.$
  • For any $t_{1},t_{2}]\subset I$ and $u\in C_{c}^{1}(N\times[t_{1},t_{2}])$,
    $\left.\int_{N}u(\cdot,t)d\mu_{t}\right|_{t=t_{1}}^{t=t_{2}}\leq\int_{t_{1}}^{t_{2}}\int_{N}(\partial_{t}u+\nabla u\cdot\vec{H}-u|\vec{H}|^{2})d\mu_{t}dt.$ (2.2)
One can quickly see that the Radon measures associated to a smooth/classical mean curvature flow, that is, a family of smooth manifolds satisfying (2.1), is a Brakke flow. A Brakke flow suffers from a serious problem: as (2.2) is only an inequality, the whole connected components of (the support of) a Brakke flow are allowed to disappear instantaneously! In the same monograph, Ilmanen [Ilmanen94_EllipReg] (see also White [White09_CurrentsVarifolds]) gives a construction via elliptic regularization of a "well-behaved" Brakke flow out of smooth initial data for which these issues don’t occur. Such a Brakke flow is known to be:
@@
Even amongst such well-behaved Brakke flows, there happens to be an issue of uniqueness, which brings us to the level set flow which originated from the viscosity method [EvansSpruck91, ChenGigaGoto91_LSF]. Later, Ilmanen [Ilmanen92_LSF] and White [White95_WSF_Top] introduced a purely set-theoretic definition using the avoidance principle. The classical avoidance principle says that two disjoint hypersurfaces in a manifold $N$ will stay disjoint, at least as long as one of them is compact. The weak set flow is defined to be a closed subset $\mathcal{M}$ of $N\times{\mathbb{R}}$ such that if a smooth mean curvature flow does not intersect $\mathcal{M}\cap\{t=t_{0}\}$, then it is disjoint from $\mathcal{M}\cap\{t=t_{1}\}$ for all $t_{1}>t_{0}$. White [White24_Avoidance] proved the following avoidance principle for weak set flows.
@@
Indeed, the support of a Brakke flow is a weak set flow, as shown in Ilmanen’s monograph. Although weak set flows may not be unique starting from a given hypersurface, there is a canonical weak set flow associated to any initial data called the level set flow: for a given set $A$ is the envelope of all families of sets initially agreeing with $A$ that satisfy the avoidance principle. As an aside, we point out that, in fact, the level set flow can even be defined this way for closed initial data that is not necessarily a regular hypersurface. While the existence and uniqueness are guaranteed, the level set flow of initial data may develop interior even if it is smooth, as first noted by Ilmanen and White [White02ICM] (and recently appeared in [IlmanenWhite24_Fattening], see also related works [AngenentIlmanenVelazquez17_fattening, LeeZhao24_MCFconical, Ketover24_Fattening]) – this fattening can be seen as an indication of nonuniqueness for level set flows. To study the level set flow even in the presence of fattening Hershkovits and White [HershkovitsWhite20_Nonfattening] introduced the notions of innermost/outermost flows. Given a closed hypersurface $\Sigma\subset{\mathbb{R}}^{3}$ that is the boundary of a domain $\Omega$, we can define:
@@
It is not immediate to define inner and outermost flows in general 33–manifolds from a 22–sided surface $M$ because such a surface need not be separating, as a cross section of $S^{2}\times S^{1}$ illustrates. In the sequel we will define such flows as the following: for a choice of normal vector $\nu$ consider the domains $\mathcal{M}^{-}$ and $\mathcal{M}^{+}$ bounded by $M$ and $M\pm\epsilon\nu$ respectively, for some small $\epsilon>0$, where $M\pm\epsilon\nu$ are surfaces that are generated by pushing $M$ in $\epsilon\nu$ direction by the exponential map. Then we define the inner and outermost flows to be the boundary component of the level set flows of $\mathcal{M}^{-}$, $\mathcal{M}^{+}$, respectively, emanating from $M$.
The third weak flow is the mean curvature flow with surgery, where one preemptively cuts out regions developing high curvature, classifies them topologically, and then continues the flow in a piecewise manner, see [Head13_MCF2convex, Lauer13_MCFSurgery, BrendleHuisken16_MCFSurgeryR3, HaslhoferKleiner17_MCFsurgery, Daniels-Holgate22_SurgeryApprox]. Locally, the high curvature regions will be modeled on round spheres or cylinders in coordinate patches, which will bound solid balls/cylinders, respectively. Regions where high curvature transitions into low curvature (i.e. "neckpinches") the surface are modeled on cylinders and at these points one "cuts" the neck at an appropriate spot and places in caps/discs of controlled geometry, in particular so that the remaining low curvature regions have curvature upper bounded by a controlled constant (which is amongst the analytical difficulties in establishing the flow with surgery). We note that the high curvature regions, after cap placement, will be homeomorphic to either $S^{n}$ or $S^{n-1}\times S^{1}$, which makes the flow with surgery a useful tool in topology. This particular consequence won’t be so relevant in this work, but the well controlled, concretely understandable change of topology across surgeries will be helpful in the sequel.
@@ 2.1. Almost regular flow and generic approximation
While these notions of weak flows seem quite different, in 33–manifolds there is a relatively complete picture to relate all of them. In [BamlerKleiner23_Multiplicity1], Bamler and Kleiner introduced a new notion of weak flow called almost regular flow, which are well behaved Brakke flows in the sense described above, along with an extra integrability condition on what they call scale functions, which measure the size of neighborhoods in which the flow is smooth. We refer the readers to [BamlerKleiner23_Multiplicity1] for the definition and properties of almost regular flows. Here, we only state properties of almost regular flows that we reference:
@@
    Generic flows, i.e., ones which only encounter mean convex singularities (see discussion and references below), are almost regular.
@@
    The innermost/outermost flows that start from a closed smooth embedded surface can be approximated by generic flows, and as a consequence, are almost regular flows for $t\geq 0$. In particular, if the level set flow from a closed smooth embedded surface is nonfattening, then it is almost regular.
@@
    (Lemma 2.7 in [BamlerKleiner23_Multiplicity1]) If $\mathcal{M}$ is almost regular, then for any test function $u\in C_{c}^{1}(N\times[t_{1},t_{2}])$ where $t_{1},t_{2}]\subset\text{Int}I$, $t\to\int_{N}u(\cdot,t)d\mu_{t}$ is continuous and the equality holds in (2.2).
While the almost regular flows have many nice properties, they still possibly develop complicated singularities that obscure the possible topological change of the flow through them. On the other hand, as we have described before, the mean curvature flow with surgery has a clear description of the topological change of the flow. Therefore, we would like to use mean curvature flow with surgery to approximate the almost regular flows we use in the sequel, particularly in showing (some of) the applications mentioned in the introduction.
In order to use mean curvature flow with surgery to approximate the almost regular flows we use in practice, we first use a well behaved (unit regular, cyclic) Brakke flow with only cylindrical and spherical singularities to approximate them. Such flows we call generic. It has been conjectured by Huisken and Angenent-Ilmanen-Chopp that mean curvature flow with generic initial data can only have cylindrical and spherical singularities, and recently, there has been various progress on this conjecture, [ColdingMinicozzi12_generic, CCMS20_GenericMCF, chodoshchoischulze2023mean, SunXue2021_initial_closed, SunXue2021_initial_conical]. Using the result of Chodosh, Choi, Mantoulidis, and Schulze [CCMS20_GenericMCF] and Chodosh, Choi, and Schulze [chodoshchoischulze2023mean], Bamler and Kleiner [BamlerKleiner23_Multiplicity1] proved that a mean curvature flow starting from a generic closed embedded surface in a three-manifold can only have cylindrical and spherical singularities. On the other hand, Daniels-Holgate [Daniels-Holgate22_SurgeryApprox] proved that the mean curvature flows with surgeries can approximate a mean curvature flow with only cylindrical and spherical singularities, where the approximation is in the Hausdorff distance. In addition, this approximation is smooth away from the singular set. Combining all the ingredients above gives the following approximation result in ${\mathbb{R}}^{3}$:
@@
Note that in the classical case, i.e., $N={\mathbb{R}}^{3}$, properly embedded hypersurfaces are automatically orientable. If $N$ is more generally simply connected, it is not hard to see this true by an intersection number argument, but we point out that it does not suffice if $N$ is merely orientable with as the example ${\mathbb{R}}P^{2}\subset{\mathbb{R}}P^{3}$ shows. With this in mind, we claim that the following holds true in general closed 33–manifolds:
@@
Let $M$ be a closed, properly embedded 22–sided surface in a closed 33–manifold NN. Then the inner and outermost flows of $M$, as defined above, can be approximated by mean curvature flows with surgeries in Hausdorff distance over any fixed time interval, and the approximation is smooth away from the singular set.
@@
The main point is to establish the existence of generic mean curvature flow ala [CCMS20_GenericMCF, chodoshchoischulze2023mean], and its subsequent approximation by surgery flows, then the statement about inner and outermost flows follows exactly as in the Euclidean case by an approximation argument as discussed in Section 7 of [BamlerKleiner23_Multiplicity1]. As is well known, in a blowup limit, singularities will still be modeled on solitons in ${\mathbb{R}}^{3}$, so we mostly just supply a sketch of this argument. Noting that because $M$ is 22–sided, one may meaningfully discuss one-sided perturbations of $M$, and the point made in the previous sentence, one readily sees from [CCMS20_GenericMCF, chodoshchoischulze2023mean] that if one encounters a nongeneric singularity of multiplicity one that the initial data can be perturbed to avoid the singularity at that spacetime point. Because $M$ is a compact surface it has finite genus and so by genus monotonicity (see Section 11 of [CCMS20_GenericMCF]) and the classification of genus 0 shrinkers by Brendle [Brendle16_genus0], only finitely many (arbitrarily small) perturbations of the initial data are required to produce a generic flow up at least to the first time a high multiplicity tangent flow develops. Now all these generic flows, as shown in Section 7 of [BamlerKleiner23_Multiplicity1], are almost regular flows; note this uses the resolution of the mean convex neighborhood conjecture by Choi, Haslhofer, and Hershkovits [ChoiHaslhoferHershkovits18_MeanConvNeighb], which the authors note carries over in the curved ambient setting. Because under rescaling as in singularity analysis the ambient metric converges smoothly to the Euclidean metric, one can see by contradiction that the resolution of the multiplicity one conjecture applies in the curved ambient setting because all the relevant geometric quantities can be arranged to evolve as they do in the flat case up to arbitrarily small errors – their argument at a high level is by showing in the case of high multiplicity convergence in a tangent flow that a certain quantity satisfies contradictory bounds so this suffices for our needs. As a result we get that for generic choice of data there will be a flow along which only mean convex singularities may develop for all time; by the resolution of the mean convex neighborhood conjecture [ChoiHaslhoferHershkovits18_MeanConvNeighb] and the barrier argument of Hershkovits and White [HershkovitsWhite20_Nonfattening, HershkovitsWhite23_Avoid], it is unique among almost regular flows starting from $M$. The approximation theorem of Daniels-Holgate also applies in this setting; most importantly, besides the mean convex neighborhood conjecture, the required local estimates for surgery also hold in the general ambient setting as shown by Haslhofer and Ketover [HaslhoferKetover19_Min2Sphere]. ∎
We next present some properties of such almost regular flows, where $M$ is, as usual, a properly embedded 22–sided surface in a given closed 33–manifold NN.
@@
Suppose that $M_{t}$ is a generic or inner/outermost flow out of $M$. Then $M_{t}$ will also be 22–sided for all regular times $t$.
@@
It’s obvious if $M_{t}$ is smooth then it will remain 22–sided, because the flow is an isotopy. In the case that $M_{t}$ has singularities, we first observe that it is easy to see for the approximating smooth flows $S_{t}$ to $M_{t}$ produced in [Daniels-Holgate22_SurgeryApprox] that $S_{t}$ will remain orientable if $M$ is. The convergence of the approximating surgery flows is smooth during regular times, giving the claim. ∎
An important subtlety to remember, though, as mentioned in the introduction, is that orientability is not necessarily inherited by the limit we discuss in the next section. In the next statement, we write $M\sim\sigma$ to indicate that $M$ is homologous to $\sigma$.
@@
Again, using the approximation by surgery flows is smooth during regular times, one only needs to note that surgery does not change the 22nd homology class of a regular surface. ∎
@@
As in the previous proofs, it suffices to note this is the case for approximating surgery flows. Now, in the surgery, components are either deleted or cylindrical regions are cut and replaced with caps, and neither of these actions do not increases the genus, giving the claim. ∎
We show in the next section that an almost regular flow will either go extinct or flow to a potentially unstable smooth minimal surface, possibly with multiplicity. If the minimal surface is unstable, we will afterwards wish to perturb the flow to ensure it doesn’t flow to it, but as discussed in the introduction, for technical reasons, it seems difficult to carry this out by just perturbing the initial data if the multiplicity is high (i.e. greater than or equal to $2$). With this in mind, we will consider the following piecewise almost regular flows in the sequel:
@@
Let $M$ be a properly embedded, 22–sided surface of a 33–manifold NN. Then a piecewise almost regular flow $M_{t}$ emanating from $M$ defined on $0,\infty)$ is given by an increasing (possibly infinite) sequence of times $t_{i}\to\infty$ for which
@@
    $M_{t}$ on $0,t_{1}]$ is an almost regular flow out of $M$ satisfying the conclusions of Proposition 2.2 above and so that $M_{t_{1}}$ is a smooth surface.
@@
    Denoting by $M_{t}^{1}$ on $0,t_{1})$ the flow from above, we let $\tilde{M}^{1}_{t_{1}}$ be a surface that is isotopic to $M^{1}_{t_{1}}$, and we define $M^{2}_{t}$ out of this perturbed surface $\tilde{M}^{1}_{t_{1}}$ as above.